# The Map Equation Goes Neural: Mapping Network Flows with Graph Neural Networks

### v3

@@ 5.2 Real-world networks with node features
We benchmark Neuromap on ten real-world datasets (Table 1) from PyTorch Geometric (PyG) [57], PyTorch Geometric Signed Directed [60], and Open Graph Benchmark (OGB) [61], and compare it against the same baselines as before. In contrast to previous works that choose a fixed number of hidden dimensions and set the maximum number of communities to a constant [34] or the “ground-truth” number of communities [36], we reflect the networks’ sizes in our choices: We set the number of hidden dimensions to $4\sqrt{n}$ and the maximum number of communities to $s=\sqrt{n}$. Our choices are based on empirical observations showing that the number of communities typically scales as $\mathcal{O}\left(\sqrt{n}\right)$ [62]. However, a few words of caution are in order: while nodes’ true communities determine the link patterns in synthetic networks, it is infeasible to obtain ground truth communities for real networks. Often, metadata labels are used as a drop-in, and the inferred communities’ quality depends on how well the metadata, which is potentially noisy, aligns with the unknown ground truth [3].
@@
For each method and architecture, we run 25 trials and report the average achieved AMI (Figure 5, a similar plot for the number of detected communities is included in Appendix B); we also report the average AMI and the detected number of communities, both with standard deviations, in tabulated form in Appendix B.
Figure 5: Average achieved AMI on real-world networks (higher is better) with $s=\sqrt{n}$. Colours indicate methods while shapes indicate neural network architectures. DiffPool ran out of memory on the ogb-arxiv dataset. Detailed tabulated results with standard deviations are included in Appendix B.
Neuromap performs best in seven cases, NOCD in three (Neuromap and NOCD achieved the same AMI on one dataset) and DiffPool in one. Remarkably, the MLP architecture performs best in four cases, GIN in three cases, GCN in two cases, and SAGE in one case. A possible explanation for the good performance of the simpler architectures could be that the loss function captures global information in the random walker’s flow patterns, making GNNs superfluous for community detection with the map equation. All methods tend to detect more communities than are present in the “ground truth”, however, this tendency is most pronounced in DMoN, Ortho, and MinCut, which may have gone unnoticed in previous evaluations where the maximum number of communities was set to a much lower, constant value [34], thus artificially preventing overfitting. Infomap is the only baseline that does not utilise node features, instead relying solely on topological information, which may explain the large number of detected communities. Comparing Neuromap’s against Infomap’s performance highlights that incorporating node features improves the detected communities’ quality substantially while drastically reducing the number of detected communities.
We repeat the same experiments with 512 hidden features and $s=\left|Y\right|$, that is, the “true” number of communities, following [36] (results in Appendix C). Limiting the number of allowed communities leads to better performance for MinCut and Ortho, however, with a few exceptions, it diminishes the performance of DMoN, NOCD, DiffPool, and Neuromap across all datasets and neural network architectures. In the case of Neuromap, imposing a lower maximum bound on the number of communities interferes with the MDL principle, limiting what models for the data may be explored.
@@ 5.3 Synthetic networks with overlapping communities
We apply Neuromap and the baselines to a small synthetic network with overlapping communities [63]. We set the maximum number of communities to $s\in\left\{2,3\right\}$, run each combination of loss function and neural network architecture for 10 trials, and keep the solutions with the lowest loss. Figure 6 shows the results obtained with GCN, Appendix D shows results for the remaining architectures.
@@ 6 Conclusion
Network science and deep learning on graphs tackle community detection from different perspectives. Community detection in network science typically relies on custom heuristic optimisation algorithms to optimise objective functions but often does not leverage recent deep learning advances. Recently, deep graph learning methods have started to incorporate methods from network science for neural graph clustering. We contribute to this young field by adapting the map equation, a popular unsupervised information-theoretic community-detection approach, as a differentiable loss function for end-to-end optimisation with (graph) neural networks through gradient descent, and use PyTorch to implement our approach, which we call Neuromap.
We evaluated Neuromap on various synthetic and real-world datasets, using different neural network architectures to detect communities. Our results show that Neuromap achieves competitive performance and detects close to the ground-truth number of communities across datasets while the baselines tend to overfit and report considerably more communities. Across all tested methods, the achieved performance depends on the used neural network architecture. However, on several real-world benchmarks, Neuromap outperforms the baselines even with simpler, non-GNN, neural network architectures. We hypothesise that this may be because the map equation builds on capturing flow patterns, which capture global information.
@@
Appendix A Results on synthetic networks with different (G)NN architectures
@@
Here, we report further results on the synthetic LFR networks with planted communities Neuromap, DMoN, NOCD, DiffPool, MinCut, and Ortho with a dense linear layer, a 2-layer MLP, a 2-layer GCN, and a 2-layer GIN. The setup is as described in the main text.
A.1 Linear-layer-based results
@@ A.4 GIN-based results
Appendix B Results on Real-World Networks
Here, we report further results on real-world networks. Specifically, we visualise the average detected number of communities per loss function and neural network architecture for each dataset. We also tabulate the average AMI and average number of detected communities, including standard deviations.
B.1 Number of detected communities on real-world networks
Figure 11: Average detected number of communities on real-world networks. Colours indicate methods while shapes indicate (G)NN architectures. The dashed horizontal lines show the correct number of communities for each dataset while the solid horizontal lines show the chosen maximum allowed number of communities. DiffPool ran out of memory on the ogb-arxiv dataset. We omit Infomap in the plot due to the large number of detected communities.
B.2 Tabulated AMI results on real-world networks
Here, we report the detailed AMI performance for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks. Table 2 lists the results on the Cora, CiteSeer, PubMed, Amazon Computer (PC), and Amazon Photo datasets. Table 3 lists the results on the Coauthor CS, Coatuhor Physics, Cora ML, Wiki CS, and ogb-arxiv datasets.
Table 2: Average AMI in % (higher is better) and their standard deviations on real-world networks. For each dataset, we highlight the best scores in bold and underline the second-best score. We tested each method with 2-layer MLP, GCN, GIN, and SAGE architectures, except for Infomap, which is not based on deep learning.
Method Arch. Cora CiteSeer PubMed PC Photo
Neuromap LIN $31.9\pm 7.9$ $19.5\pm 1.6$ $17.2\pm 12.2$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
MLP $36.6\pm 2.4$ $15.8\pm 1.1$ $23.2\pm 1.8$ $22.4\pm 8.6$ $39.8\pm 11.2$
GCN $38.5\pm 2.0$ $16.2\pm 1.3$ $20.8\pm 1.2$ $\textbf{52.1}\pm 2.3$ $\textbf{65.0}\pm 3.8$
GIN $\textbf{46.1}\pm 2.0$ $23.1\pm 1.2$ $\textbf{24.6}\pm 1.7$ $43.2\pm 2.4$ $58.2\pm 6.0$
SAGE $41.6\pm 1.9$ $18.3\pm 1.2$ $20.4\pm 1.0$ $43.6\pm 4.8$ $48.1\pm 5.8$
DMoN LIN $3.8\pm 1.1$ $2.9\pm 1.3$ $0.4\pm 0.7$ $2.7\pm 1.9$ $5.9\pm 3.5$
MLP $8.0\pm 1.2$ $6.5\pm 0.9$ $13.9\pm 0.6$ $19.6\pm 3.3$ $34.2\pm 2.3$
GCN $32.8\pm 1.2$ $16.3\pm 0.8$ $17.1\pm 0.3$ $42.1\pm 0.5$ $48.7\pm 0.5$
GIN $33.8\pm 1.0$ $19.3\pm 1.3$ $16.3\pm 0.5$ $28.5\pm 1.8$ $43.6\pm 1.1$
SAGE $32.4\pm 1.3$ $16.0\pm 1.0$ $15.4\pm 0.2$ $32.5\pm 1.4$ $43.2\pm 0.7$
NOCD LIN $5.1\pm 1.9$ $2.4\pm 1.3$ $3.7\pm 1.7$ $1.2\pm 2.7$ $0.1\pm 0.3$
MLP $\textbf{46.1}\pm 1.1$ $\textbf{34.8}\pm 0.9$ $20.5\pm 2.4$ $45.5\pm 2.5$ $60.5\pm 3.4$
GCN $41.1\pm 0.9$ $20.2\pm 0.8$ $21.2\pm 0.6$ $\underline{51.9}\pm 1.9$ $63.2\pm 0.9$
GIN $43.4\pm 1.1$ $\underline{26.4}\pm 1.2$ $21.3\pm 0.6$ $34.2\pm 1.9$ $55.3\pm 1.9$
SAGE $40.7\pm 0.9$ $19.8\pm 0.8$ $19.6\pm 0.4$ $51.1\pm 1.6$ $\underline{63.7}\pm 1.6$
DiffPool LIN $0.4\pm 1.1$ $1.1\pm 1.9$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
MLP $9.0\pm 1.2$ $6.7\pm 1.0$ $18.9\pm 1.4$ $1.6\pm 0.4$ $3.1\pm 0.9$
GCN $39.8\pm 2.0$ $18.2\pm 1.3$ $\underline{23.7}\pm 1.5$ $37.0\pm 8.0$ $53.8\pm 7.5$
GIN $30.9\pm 3.6$ $21.3\pm 2.3$ $16.3\pm 3.9$ $14.5\pm 3.2$ $32.8\pm 4.6$
SAGE $38.3\pm 2.4$ $16.4\pm 1.4$ $21.9\pm 1.5$ $3.1\pm 0.3$ $16.7\pm 6.1$
MinCut LIN $26.1\pm 2.5$ $16.5\pm 1.3$ $2.2\pm 1.0$ $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
MLP $34.7\pm 6.6$ $20.5\pm 0.9$ $5.1\pm 2.4$ $0.0\pm 0$ 0.0 $5.2\pm 9.9$
GCN $26.2\pm 1.5$ $12.0\pm 0.7$ $11.0\pm 2.8$ $15.9\pm 21.2$ $45.2\pm 13.6$
GIN $39.3\pm 3.8$ $18.9\pm 0.7$ $19.8\pm 4.3$ $6.8\pm 6.8$ $35.5\pm 19.3$
SAGE $31.5\pm 1.5$ $16.2\pm 0.8$ $15.5\pm 0.3$ $41.6\pm 0.6$ $46.9\pm 0.5$
Ortho LIN $7.0\pm 0.7$ $5.6\pm 0.7$ $16.5\pm 0.4$ $18.8\pm 9.0$ $24.7\pm 11.4$
MLP $5.3\pm 0.5$ $3.8\pm 0.4$ $10.6\pm 0.4$ $19.8\pm 5.5$ $30.6\pm 4.0$
GCN $22.9\pm 0.9$ $10.9\pm 0.7$ $13.8\pm 0.5$ $40.7\pm 0.9$ $46.0\pm 1.4$
GIN $30.7\pm 1.3$ $17.8\pm 0.9$ $13.6\pm 0.5$ $34.0\pm 0.9$ $42.3\pm 0.7$
SAGE $18.0\pm 0.9$ $8.4\pm 0.8$ $9.9\pm 0.4$ $35.4\pm 3.9$ $46.3\pm 1.6$
Infomap $35.2\pm 0.2$ $23.6\pm 0.1$ $16.0\pm 0.1$ $49.5\pm 0.5$ $57.3\pm 1.1$

### v4

@@ 5.2 Real-world networks with node features
We benchmark Neuromap on ten real-world datasets (Table 1) from PyTorch Geometric [59], PyTorch Geometric Signed Directed [62], and Open Graph Benchmark [63], and compare it against the same baselines as before. In contrast to previous works that choose a fixed number of hidden dimensions and set the maximum number of communities to a constant [36] or the “ground-truth” number of communities [38], we reflect the networks’ sizes in our choices: We set the number of hidden dimensions to $4\sqrt{n}$ and the maximum number of communities to $s=\sqrt{n}$. Our choices are based on empirical observations showing that the number of communities typically scales as $\mathcal{O}\left(\sqrt{n}\right)$ [64]. However, a few words of caution are in order: while nodes’ true communities determine the link patterns in synthetic networks, it is generally infeasible to obtain ground truth communities for real networks. Often, metadata labels are used as a drop-in, and the inferred communities’ quality depends on how well the metadata, which is potentially noisy, aligns with the unknown ground truth [3]. Moreover, determining the number of communities in a network is hard and setting $s=\sqrt{n}$ should be seen as a simplification rather than an attempt to guess the exact number.
@@
For each method and architecture, we run 25 trials and show the average achieved AMI in Figure 5; we include a similar plot for the number of detected communities as well as the average AMI and the detected number of communities, both with standard deviations, in tabulated form in Appendix C. When several of the best-performing methods achieve similar AMI, we use an independent two-sample t-test to determine whether one of them can be considered to perform better than the other (see Appendix C). In cases where their performances do not differ significantly, we mark both as best.
Figure 5: Average achieved AMI on real-world networks (higher is better) with $s=\sqrt{n}$. Colours indicate methods while shapes indicate neural network architectures. DiffPool ran out of memory on the ogb-arxiv dataset. Detailed tabulated results with standard deviations are included in Appendix C.
Neuromap and NOCD are amongst the best performers in seven cases and DiffPool in two. The GCN architecture performs best in seven cases, MLP and SAGE in four cases, and GIN in two cases. A possible explanation for why the simpler linear layer and MLP architectures perform well in several cases could be that the map equation loss function captures global information in the random walker’s flow patterns, making GNNs superfluous in some cases. All methods tend to detect more communities than are present in the “ground truth”, however, this tendency is most pronounced in DMoN, Ortho, and MinCut, which may have gone unnoticed in previous evaluations where the maximum number of communities was set to a much lower, constant value [36], thus artificially preventing overfitting. Infomap is the only baseline that does not utilise node features; instead, it relies solely on topological information, which may explain the large number of detected communities. Comparing Neuromap’s performance against Infomap’s performance suggests that incorporating node features substantially improves the detected communities’ quality in most cases (see Appendix C for significance tests) while drastically reducing the number of detected communities.
We repeat the same experiments with 512 hidden features and $s=\left|Y\right|$, that is, the “true” number of communities, following [38] (results in Appendix D). Limiting the number of allowed communities often leads to better performance for DMoN, MinCut, and Ortho, however, with a few exceptions, it diminishes the performance of Neuromap, NOCD, and DiffPool across all datasets and neural network architectures. Appendix E tabulates the differences in average AMI score between setting the hidden features to $4\sqrt{n}$ and $s=\sqrt{n}$ versus using $512$ hidden features and $s=\left|Y\right|$. In the case of Neuromap, imposing a lower bound on the number of communities interferes with the MDL principle, limiting what models for the data may be explored.
@@ 5.3 Synthetic networks with overlapping communities
We apply Neuromap and the baselines to a small synthetic network with overlapping communities [65]. We set the maximum number of communities to $s\in\left\{2,3\right\}$, run each combination of loss function and neural network architecture for 10 trials, and keep the solutions with the lowest loss. Figure 6 shows the results obtained with GCN, Appendix F shows results for the remaining architectures.
@@ 6 Conclusion
Network science and deep learning on graphs tackle community detection from different perspectives. Community detection in network science typically relies on custom heuristic optimisation algorithms to optimise objective functions but often does not leverage recent deep learning advances. Recently, deep graph learning methods have started to incorporate methods from network science for deep graph clustering. We contribute to this young field by adapting the map equation, a popular unsupervised information-theoretic community-detection approach, as a differentiable loss function for end-to-end optimisation with (graph) neural networks through gradient descent, and use PyTorch to implement our approach, which we call Neuromap.
We evaluated Neuromap on various synthetic and real-world datasets, using different neural network architectures to detect communities. Our results show that Neuromap achieves competitive performance and detects close to the ground-truth number of communities across datasets while the baselines tend to overfit and report considerably more communities. Across all tested methods, the achieved performance depends on the used neural network architecture. However, on several real-world benchmarks, Neuromap outperforms several of the the baselines even with simpler, non-GNN, neural network architectures. We hypothesise that this may be because the map equation builds on capturing flow patterns, which contain global information.
@@
Appendix A Map equation details
Let $G=\left(V,E\right)$ be a graph with nodes $V$, links $E$, and let $w_{uv}\in\mathbb{R}^{+}_{0}$ denote the non-negative weight on the link from node $u$ to $v$. Further, let $p_{u}$ be node $u$’s ergodic visit rate, which can be calculated in closed form as $p_{u}=s_{u}/\sum_{v\in V}s_{v}$ in undirected graphs, where $s_{u}=\sum_{v\in V}w_{uv}$ is node $u$’s strength. In directed graphs, $p_{u}$ can be calculated numerically with smart teleportation and a power iteration by solving the recursive set of equations $p_{v}=\alpha\frac{s_{v}^{\text{out}}}{\sum_{u\in V}s_{u}^{\text{out}}}+\left(1-\alpha\right)\sum_{u\in V}p_{u}t_{uv}$, where $s_{u}^{\text{out}}=\sum_{v\in V}w_{uv}$ is node $u$’s out-strength, $t_{uv}=w_{uv}/\sum_{v\in V}w_{uv}$ is the random walker’s transition probability from node $u$ to node $v$, and $\alpha$ is a teleportation parameter, typically set to $\alpha=0.15$ [45]. Smart teleportation is similar to PageRank [67] but, instead of uniformly teleporting to nodes, the random walker teleports to links proportional to their weight.
For a given partition $\mathsf{M}$ of the nodes into modules, the map equation [16] calculates the average per-step description length – also called codelength – for describing the position of a random walker on the graph:
$L\left(\mathsf{M}\right)=qH\!\left(Q\right)+\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}H\!\left(P_{\mathsf{m}}\right),$ (4)
Here, $q=\sum_{\mathsf{m}}q_{\mathsf{m}}$ is the random walker’s module entry rate, $q_{\mathsf{m}}=\sum_{u\notin\mathsf{m}}\sum_{v\in\mathsf{m}}p_{u}t_{uv}$ is module $\mathsf{m}$’s entry rate, and $Q=\left\{q_{\mathsf{m}}/q\mid\mathsf{m}\in\mathsf{M}\right\}$ is the set of normalised module entry rates; $p_{\mathsf{m}}=\mathsf{m}_{\text{exit}}+\sum_{u\in\mathsf{m}}p_{u}$ is the rate at which the random walker moves in module $\mathsf{m}$, including the exit rate $\mathsf{m}_{\text{exit}}=\sum_{u\in\mathsf{m}}\sum_{v\notin\mathsf{m}}p_{u}t_{uv}$, and $P_{\mathsf{m}}=\left\{\mathsf{m}_{\text{exit}}/p_{\mathsf{m}}\right\}\cup\left\{p_{u}/p_{\mathsf{m}}\mid u\in\mathsf{m}\right\}$ is the set of normalised node visit rates and exit rates for module $\mathsf{m}$.
A.1 Rewriting the map equation
The map equation can be rewritten by expanding the entropy terms and cancelling common factors
$L\left(\mathsf{M}\right)$ $=qH\!\left(Q\right)+\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}H\!\left(P_{\mathsf{m}}\right)$ (5)
[…]
Applying logarithm rules gives
$=-\!\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q_{\mathsf{m}}+\!\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q$ (7)
$\qquad-\!\sum_{\mathsf{m}\in\mathsf{M}}\mathsf{m}_{\text{exit}}\log_{2}\mathsf{m}_{\text{exit}}+\!\sum_{\mathsf{m}\in\mathsf{M}}\mathsf{m}_{\text{exit}}\log_{2}p_{\mathsf{m}}-\!\sum_{\mathsf{m}\in\mathsf{M}}\sum_{u\in\mathsf{m}}p_{u}\log_{2}p_{u}+\!\sum_{\mathsf{m}\in\mathsf{M}}\sum_{u\in\mathsf{m}}p_{u}\log_{2}p_{\mathsf{m}},$ (8)
and further simplification yields
$=q\log_{2}q-\!\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q_{\mathsf{m}}+\!\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}\log_{2}p_{\mathsf{m}}-\!\sum_{\mathsf{m}\in\mathsf{M}}\mathsf{m}_{\text{exit}}\log_{2}\mathsf{m}_{\text{exit}}-\!\sum_{u\in V}p_{u}\log_{2}p_{u}.$ (9)
In undirected networks where $q_{\mathsf{m}}=\mathsf{m}_{\text{exit}}$, this can be further simplified [20] to
$=q\log_{2}q-2\sum_{\mathsf{m}\in\mathsf{M}}q_{\mathsf{m}}\log_{2}q_{\mathsf{m}}+\!\sum_{\mathsf{m}\in\mathsf{M}}p_{\mathsf{m}}\log_{2}p_{\mathsf{m}}-\!\sum_{u\in V}p_{u}\log_{2}p_{u}.$ (10)
The last term in Equation 10, that is, the nodes’ contribution to the codelength, is constant because it does not depend on the module structure, and can be omitted during optimisation.
A.2 Node flow with soft cluster assignments
While the nodes’ codelength contribution is constant with hard clusters (Equation 10), expressing the map equation with soft cluster assignments allows modelling the flow of nodes that are assigned to more than one module in at least two different ways which we illustrate in Figure 7.
The first option is to reflect the nodes’ partial cluster assignment for assigning codewords. Consider node $u$ with visit rate $p_{u}$ and partial assignments of $\frac{1}{2}$ to clusters $\mathsf{m}_{1}$ whose usage rate is $p_{\mathsf{m}_{1}}$ and $\frac{1}{2}$ to cluster $\mathsf{m}_{2}$ whose usage rate is $p_{\mathsf{m}_{2}}$. Then, half of $u$’s flow, that is $\frac{p_{u}}{2}$, falls into each of the two clusters and $u$’s contribution to the overall codelength is $-\frac{p_{u}}{2}\log_{2}\frac{p_{u}}{2p_{\mathsf{m}_{1}}}-\frac{p_{u}}{2}\log_{2}\frac{p_{u}}{p_{2\mathsf{m}_{2}}}$. In general, let node $u$’s assignment to cluster $\mathsf{m}_{i}$ be $s_{ui}\in\left[0,1\right]$ with $\sum_{i}s_{ui}=1$, and let $\mathsf{m}_{i}$’s usage rate be $p_{\mathsf{m}_{i}}$. Then, node $u$’s contribution to the overall codelength is $-\sum_{i}s_{ui}p_{u}\log_{2}\frac{s_{ui}p_{u}}{p_{\mathsf{m}_{i}}}$. Essentially, this approach splits each node with assignments to more than one cluster into its corresponding parts, resulting in potentially many small pieces with low visit rates.
The second option is to treat the nodes as indivisible when computing their codeword lengths but reflecting their soft assignments in the contribution to the overall codelength. Using the same example as above, node $u$’s contribution to the overall codelength is $-\frac{p_{u}}{2}\log_{2}\frac{p_{u}}{p_{\mathsf{m}_{1}}}-\frac{p_{u}}{2}\log_{2}\frac{p_{u}}{p_{\mathsf{m}_{2}}}$, which is lower because the values inside the logarithms are bigger. In general, $u$’s contribution to the codelength is $-\sum_{i}s_{ui}p_{u}\log_{2}\frac{p_{u}}{p_{\mathsf{m}_{i}}}$, where the difference to the previous option is that $s_{ui}$ does not appear inside the logarithm.
[…]
Figure 7: Modelling node flow with soft cluster assignments. The labels show each node’s contribution to the overall codelength based on their visit rates. (a) The one-level partition where all nodes belong to the same community. (b) With hard communities, the middle node must be assigned to either the blue or orange community. Here, this leads to a higher codelength than for the one-level partition. (c) Reflecting nodes’ partial assignments in the codelength contributions means splitting them into several smaller pieces whose visit rates sum to the original node’s visit rate. Important objects are split into several less important objects, requiring longer codewords because of their lower visit rates. Overall, this tends to increase the codelength and can prevent identifying overlapping modules. (d) Treating nodes’ contribution to the codelength as constant keeps the nodes intact and allows important nodes to retain their higher visit rates, leading to shorter codewords. As the codelength highlights, only this approach would identify communities in this example.
Appendix B Results on synthetic networks with different (G)NN architectures
@@
Here, we report further results on the synthetic LFR networks with planted communities for Neuromap, DMoN, NOCD, DiffPool, MinCut, and Ortho using the following architectures: a dense linear layer, a 2-layer MLP, a 2-layer GCN, and a 2-layer GIN. The setup is as described in the main text.
B.1 Linear-layer-based results
@@ B.4 GIN-based results
Appendix C Results on Real-World Networks
Here, we report further results on real-world networks for setting the number of hidden dimensions to $4\sqrt{n}$ and the maximum number of communities to $s=\sqrt{n}$. Specifically, we visualise the average detected number of communities per loss function and neural network architecture for each dataset. We tabulate the average AMI and average number of detected communities, including standard deviations. Two-sample t-tests show when Neuromap or a baseline performs significantly better than the other.
C.1 Number of detected communities on real-world networks
Figure 12: Average detected number of communities on real-world networks. Colours indicate methods while shapes indicate (G)NN architectures. The dashed horizontal lines show the correct number of communities for each dataset while the solid horizontal lines show the chosen maximum allowed number of communities, here $s=\sqrt{n}$. DiffPool ran out of memory on the ogb-arxiv dataset. We omit Infomap in the plot due to the large number of detected communities.
C.2 Tabulated AMI results on real-world networks
Here, we report the detailed AMI performance for Neuromap, DMoN, NOCD, DiffPool, MinCut, Ortho, and Infomap on real-world networks for setting the number of hidden dimensions to $4\sqrt{n}$ and the maximum number of communities to $s=\sqrt{n}$. When several of the best-performing methods achieve similar AMI, we use an independent two-sample t-test to determine whether one of them can be considered to perform better than the other. In cases where their performances do not differ significantly, we mark both as best.